# CryoFM: A Flow-based Foundation Model for Cryo-EM Densities

### v1

@@ 5.5 Ablation & Discussion
In this section, we perform ablation studies to analyze the impact of different design choices and hyper-parameters. We briefly present some conclusions here and refer the reader to Appendix F for details.
Figure 8: MSE loss on test set as a function of total training compute. The estimation of training compute is consistent with Peebles & Xie (2023a).
@@
Figure 10: Averaged ranking ($\downarrow$) across downstream tasks for different hyperparameters.
Figure 11: Unconditional sampling from models in different spaces. For latent flow models (LFM), the high-frequency information contains more noise than signal.
@@
Data normalization enhances downstream performance across all tasks. Fig. 9 demonstrates that models trained with normalized data consistently improves the $\text{FSC}_{\text{AUC}}$ metrics compared to the one without normalization in three downstream tasks. The observation is in line with some DDPM-based work (Yim et al., 2023b; Rombach et al., 2022), as a significant change in the variance of the data can make the model difficult to train (Karras et al., 2022).
@@
Flow models trained in the voxel space converge much faster and better than the latent flow models. We conducted experiments on training the model in the latent space (Rombach et al., 2022). Fig. 11 illustrates the unconditional sampling results from two latent flow models (LFM), where the high-frequency information contains more noise than signal. The observation concides with that of Crowson et al. (2024), where they found that latent diffusion models fail to generate fine details. We suspect that the VAE is under-trained due to a small amount of data, and leave for future work.
@@ 6 Conclusion
In this study, we present CryoFM, a flow matching-based foundation model that learns the prior distribution of high-quality cryo-EM densities. During inference, we derive the degradation operators for specific tasks, allowing protein densities to be sampled from the posterior distribution based on given observations. CryoFM demonstrates versatility by restoring protein densities across four distinct tasks without fine-tuning, showcasing the potential of deep generative models as a powerful prior in cryo-EM. While promising, CryoFM has limitations: though we use real data with synthetic noise for comparison, applying the method directly to real-world noisy densities remains challenging. Additionally, our work does not address reconstructing 3D densities from raw 2D particles, but we believe CryoFM can contribute to solving these complex tasks in future work.
Appendix A Additional Related Work
A.1 Density modification and denoising in cryo-EM
@@
Traditional density modification methods often apply Fourier space weighting, utilizing frequency-dependent scaling to suppress noise and enhance signal (Jakobi et al., 2017; Ramírez-Aportela et al., 2020; Terwilliger et al., 2020a; Kaur et al., 2021). These heuristic-based approaches, like Wiener-style deblurring (Ramírez-Aportela et al., 2020), tend to rely on local resolution estimates and filtering strategies but may struggle with intricate structural details due to limited priors. On the deep learning side, pretrained models like Blush (Kimanius et al., 2024) further refine density maps by denoising half-maps during iterative refinement. However, they can introduce hallucinated details in lower-resolution maps. In contrast, methods such as M (Tegunov et al., 2021) and spIsoNet (Liu et al., 2024) avoid external data, with M leveraging noise2noise (Moran et al., 2020) for denoising independent half-maps, and spIsoNet addressing anisotropic signal distributions through self-supervised learning. Though slower, these approaches tend to reduce the risk of hallucinations and offer better robustness, but their power is often limited by not being data-driven.
A.2 Missing wedge problem in cryo-ET
@@
A.3 Ab initio modeling in cryo-EM
@@
A.4 Diffusion/Flow-based Model
Flow-based models (Lipman et al., 2022; Liu et al., 2022a) and denoising diffusion probablistic models (DDPM) (Sohl-Dickstein et al., 2015; Ho et al., 2020) are two modern deep generative models. They exhibit many similarities in technical details. DDPM implements an iterative refinement process by learning to gradually denoise a sample from a normal distribution. It has achieved the state-of-the-art results on many generative tasks, including image generation (Rombach et al., 2022; Podell et al., 2023; Peebles & Xie, 2023b), video generation (Blattmann et al., 2023), and molecule generation (Yim et al., 2023b; Abramson et al., 2024; Ingraham et al., 2023; Watson et al., 2023; Wang et al., 2024), etc. Recently, diffusion models have been adopted in the cryo-EM field. Kreis et al. (2022) traverses the latent space of cryoDRGN (Zhong et al., 2021a) with a diffusion model, while Wang et al. (2024) refines structures for model building by iteratively denoising the density. Flow-based models regress a vector field that generate a disired probability path. Their simple and efficient implementation enables fast learning. These models have shown success in various domains, including image generation (Esser et al., 2024) and moleculer generation (Yim et al., 2023b; Bose et al., 2023). The DDPM objective can also be unified into flow-based models by converting it into a probability flow ODE (Song et al., 2021b).
A.5 Vision Transformers for Diffusion Models
Diffusion transformers (Peebles & Xie, 2023a) have demonstrated significant scalability and generative capabilities in image-related tasks (Esser et al., 2024; Hoogeboom et al., 2023; Hatamizadeh et al., 2024; Zhou et al., 2024). In particular, HDiT (Crowson et al., 2024) leverages the inherent hierarchical nature of visual patterns in its model design. By integrating the characteristics of Diffusion Transformer (Peebles & Xie, 2023a) and Hourglass transformers (Nawrot et al., 2022), and employing local attention mechanisms (Hassani et al., 2023), HDiT offers an efficient model structure suitable for training the diffusion model in the data space.
Appendix B Additional Detailed Information of CryoFM
B.1 Implementation Details
@@
B.2 Likelihood Estimation
Given a data point ${\mathbf{x}}_{0}\in\mathbb{R}^{n}$, we estimate its likelihood by solving a probability flow ODE (Song et al., 2021b; Lipman et al., 2022; Chen et al., 2018). We start with the continuity equation of a velocity field $v_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$:
$\frac{d}{dt}\log p_{t}({\mathbf{x}}_{t})+\nabla\cdot v_{t}({\mathbf{x}}_{t})=0.$
@@ B.2 Likelihood Estimation
$\frac{d}{dt}\left[\begin{array}[]{cc}{\mathbf{x}}_{t}\\ f_{t}\end{array}\right]=\left[\begin{array}[]{cc}v_{t}({\mathbf{x}}_{t})\\ \nabla\cdot v_{t}({\mathbf{x}}_{t})\end{array}\right],$
@@
$\frac{d}{dt}\left[\begin{array}[]{cc}{\mathbf{x}}_{t}\\ f_{t}\end{array}\right]=\left[\begin{array}[]{cc}v_{t}({\mathbf{x}}_{t})\\ {\mathbf{z}}^{T}\nabla v_{t}({\mathbf{x}}_{t}){\mathbf{z}}\end{array}\right],$
@@
Appendix C Additional information of the experiments
C.1 Metrics
@@
C.2 Spectral Noise Denoising
@@
Appendix D Estimation of the degradation operators
D.1 Spectral noise power estimation
@@
D.2 Anisotropic noise power estimation
@@
where $h(\nu)$ is the average noise power derived from the the half-map GSFSC in Section D.1.
@@
Appendix E Additional Algorithms
@@
E.1 Missing Wedge Restoration
For certain tasks, such as missing wedge restoration, it is not necessary to compute the gradient with respect to the loss function between the observation ${\mathbf{y}}$ and the degraded signal $\mathcal{A}{\mathbf{x}}$, since we can directly attain the optimum in an analytical form. The missing wedge restoration is an in-painting task, where a noiseless part of the signal is already observed. $\mathcal{A}$ is an operator extracting a part of the signal, hence the inverse of $\mathcal{A}$ just fills the observed part into the signal. In Alg. 2, we formulate it as the pseudo-inverse operator $\mathcal{A}^{\dagger}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ (Kawar et al., 2021; 2022).
@@ E.1 Missing Wedge Restoration
a pretrained vector field $v_{\theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, a degradation operator $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, an obseravation ${\mathbf{y}}\in\mathbb{R}^{m}$, number of steps $N$
@@
for $t\in[1,1-\frac{1}{N},1-\frac{2}{N},\cdots,0]$ do
     $\hat{{\mathbf{x}}}_{0}\leftarrow{\mathbf{x}}_{t}-t\cdot v_{\theta}(t,{\mathbf{x}}_{t})$
@@
     ${\mathbf{x}}_{t-1}\leftarrow\frac{Nt-1}{Nt}\cdot{\mathbf{x}}_{t}+\frac{1}{Nt}\cdot\hat{{\mathbf{x}}}^{\prime}_{0}$
@@
E.2 Ab initio Modeling
For ab initio modeling, the $k$-th degraded volume (we will name it the projection later) is $\mathcal{A}^{(k)}({\bm{V}})=\mathcal{P}(\phi^{(k)},{\bm{V}})$, where $\phi^{(k)}\in\mathbb{SO}(3)\times\mathbb{R}^{2}$ is the pose. The posterior sampling of ab initio modeling is more complex than the previous tasks in that the pose $\phi^{(k)}$ can not be easily pre-determined. The setting is closely related to blind diffusion posterior sampling (Chung et al., 2023a; Kapon et al., 2024; Laroche et al., 2024; Gan et al., 2024), where the degradation operator is unknown. We find $\phi^{(k)}$ in the sampling process, by minimizing the discrepancy between the projection and the observation ${\bm{y}}^{(k)}$ using a coarse-to-fine strategy (Zhong et al., 2021b). The discrepancy is defined by the negative correlation between a projection ${\mathbf{p}}$ and an observation ${\mathbf{y}}$:

### v2

@@ 5.5 Ablation & Discussion
In this section, we perform ablation studies to analyze the impact of different design choices and hyper-parameters. We briefly present some conclusions here and refer the reader to Appendix H.
Figure 8: MSE loss on test set as a function of total training compute. The estimation of training compute is consistent with Peebles & Xie (2023).
@@
Figure 10: Averaged ranking ($\downarrow$) across downstream tasks for hyperparameters.
Figure 11: Unconditional sampling from models in different spaces. For latent flow models (LFM), the high-frequency information contains more noise.
@@
Data normalization enhances downstream performance across all tasks. Data normalization is the last step in data processing (as referenced in Section C.2). A density with a value ranging from $[0,1]$ is subtracted by $0.04$ and then divided by $0.09$. Fig. 9 demonstrates that models trained with normalized data consistently improves the $\text{FSC}_{\text{AUC}}$ metrics compared to the one without normalization in three downstream tasks. The observation is in line with some DDPM-based work (Yim et al., 2023b; Rombach et al., 2022), as a significant change in the variance of the data can make the model difficult to train (Karras et al., 2022).
@@
Flow models trained in the voxel space converge much faster and better than the latent flow models. We conducted experiments on training the model in the latent space (Rombach et al., 2022). Fig. 11 illustrates the unconditional sampling results from two latent flow models (LFM), where the high-frequency information is more noisy. The observation concides with that of Crowson et al. (2024), where they found that latent diffusion models fail to generate fine details. We suspect that the VAE is under-trained due to a small amount of data, and we will leave this for future work.
@@ 6 Conclusion
In this study, we present CryoFM, a flow matching-based foundation model that learns the prior distribution of high-quality cryo-EM densities. During inference, we derive the forward operators for specific tasks, allowing protein densities to be sampled from the posterior distribution based on given observations. CryoFM demonstrates versatility by restoring protein densities across four distinct tasks without fine-tuning, showcasing the potential of deep generative models as a powerful prior in cryo-EM. While promising, CryoFM has limitations: though we use real data with synthetic noise for comparison, applying the method directly to real-world noisy densities remains challenging. Additionally, our work does not address reconstructing 3D densities from raw 2D particles, but we believe CryoFM can contribute to solving these complex tasks in future work.
Appendix A Discussion on Flow Matching
A.1 Concepts in Flow Matching
Fig. 12 illustrates three main concepts in flow matching. Here, the horizontal bar represents the time dimension, and each vertical slice represents a time-dependent probability distribution with only one dimension.
Figure 12: An illustration of three main concepts in flow matching: (i) the vector field ${\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$, (ii) the probability density path $p_{t}({\mathbf{x}}_{t})$, and (iii) the flow of ${\mathbf{x}}_{t}$. The approximated posterior mean $\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})$ in Alg. 1 is also depicted on the right part.
A.2 Choice of Flow Matching for Foundation Models
In this section, we discuss the choice of employing flow matching as the backbone of the cryo-EM foundation model. The main motivation is that (i) generative modeling is a trend in building modern foundation models, and (ii) flow matching excels at generative modeling.
Generative modeling is a trend in building modern foundation models. ChatGPT (Achiam et al., 2023) has disrupted the entire NLP (natural language processing) area with its extremely strong generative power. It brings about a paradigm shift from representation learning to generative modeling. For representation learning, one of the most prominent models is BERT (Devlin, 2018), which performs well in language understanding but is not adept at performing generation. Similar trends can also be observed in computational biology. Two typical cases in this field are ESM-3 and AlphaFold3. Built upon ESM (Rives et al., 2021) (a BERT-style model), Hayes et al. (2024) developed ESM-3 that has the ability to generate protein sequences. A major change from AlphaFold2 (Jumper et al., 2021) to AlphaFold3 (Abramson et al., 2024) is the diffusion module that can be utilized to generate an ensemble of structures.
Flow matching excels at generative modeling. Diffusion-based foundation models have had a significant impact on generative modeling in AI for biology. AlphaFold3 (Abramson et al., 2024), RFDiffusion (Watson et al., 2023), and Chroma (Ingraham et al., 2023) are all based on diffusion models (DDPM). ESM-3 (Hayes et al., 2024), although not explicitly stated in its manuscript, is closely related to discrete diffusion models. One of the crucial reasons for the success of DDPM resides in iterative refinement, which gradually refines the generated samples to a more superior state. The iterative nature of DDPM enables it to outperform its counterparts with one-step generation methods, such as VAE and GAN. Flow matching can be regarded as a variant of DDPM and shares many similar properties with DDPM. We prefer flow matching over DDPM since it is easier to train and can converge faster. In the early stage of this study, we trained both DDPM and flow matching. We found that flow matching can yield similar results while incurring less computational cost.
Appendix B Additioiant onal Related Work
B.1 Density modification and denoising in cryo-EM
@@
Traditional density modification methods often apply Fourier space weighting, utilizing frequency-dependent scaling to suppress noise and enhance signal (Jakobi et al., 2017; Ramírez-Aportela et al., 2020; Terwilliger et al., 2020a; Kaur et al., 2021). These heuristic-based approaches, like Wiener-style deblurring (Ramírez-Aportela et al., 2020), tend to rely on local resolution estimates and filtering strategies but may struggle with intricate structural details due to limited priors. On the deep learning side, Blush (Kimanius et al., 2024) further refine density maps by denoising half-maps during iterative refinement. However, they can introduce hallucinated details in lower-resolution maps. In contrast, methods such as M (Tegunov et al., 2021) and spIsoNet (Liu et al., 2024) avoid external data, with M leveraging noise2noise (Moran et al., 2020) for denoising independent half-maps, and spIsoNet addressing anisotropic signal distributions through self-supervised learning. Though slower, these approaches tend to reduce the risk of hallucinations and offer better robustness, but their power is often limited by not being data-driven. In this paper, we selected DeepEMhancer (Sanchez-Garcia et al., 2021), EMReady (He et al., 2023) and spIsoNet as the baselines for the denoising tasks.
DeepEMhancer DeepEMhancer is a deep learning-based tool designed to enhance cryo-EM density maps, improving their interpretability and aiding in structural analysis. The method employs a convolutional neural network (CNN) trained on pairs of experimental cryo-EM maps and corresponding locally sharpened maps. This training enables DeepEMhancer to learn the mapping between low-quality input maps and their high-quality counterparts. The pretrained model processes new cryo-EM maps to produce enhanced versions with improved clarity and detail, facilitating more accurate structural interpretations.
EMReady Similar to DeepEMhancer, EMReady is a computational method designed to enhance cryo-EM density maps by simultaneously applying local and non-local denoising techniques. The method uses a Swin-Conv-3DUNet and trained on pairs of experimental cryo-EM maps and correponding synthetic maps from the atomic models. The pretrained model can enhance cryo-EM maps for better interpretations.
spIsoNet spIsoNet is a self-supervised deep learning method developed to address the preferred orientation problem in single-particle cryo-EM. This issue arises when particles predominantly adopt specific orientations during imaging, leading to anisotropic data and potential inaccuracies in 3D reconstructions. The method adopts a U-Net architecture and leverages the half maps in cryo-EM reconstruction to learn representations from the overrepresented orientations, recovering information for the underrepresented or missing views.
B.2 Missing wedge problem in cryo-ET
@@
B.3 Ab initio modeling in cryo-EM
@@
B.4 Diffusion/Flow-based Model
Flow-based models (Lipman et al., 2022; Liu et al., 2022a) and denoising diffusion probablistic models (DDPM) (Sohl-Dickstein et al., 2015; Ho et al., 2020) are two modern deep generative models. They exhibit many similarities in technical details. DDPM implements an iterative refinement process by learning to gradually denoise a sample from a normal distribution. It has achieved the state-of-the-art results on many generative tasks, including image generation (Rombach et al., 2022; Podell et al., 2023; Peebles & Xie, 2023), video generation (Blattmann et al., 2023), and molecule generation (Yim et al., 2023b; Abramson et al., 2024; Ingraham et al., 2023; Watson et al., 2023; Wang et al., 2024b), etc. Recently, diffusion models have been adopted in the cryo-EM field. Kreis et al. (2022) traverses the latent space of cryoDRGN (Zhong et al., 2021a) with a diffusion model, while Wang et al. (2024a) refines structures for model building by iteratively denoising the density. Flow-based models regress a vector field that generate a disired probability path. Their simple and efficient implementation enables fast learning. These models have shown success in various domains, including image generation (Esser et al., 2024) and moleculer generation (Yim et al., 2023b; Bose et al., 2023). The DDPM objective can also be unified into flow-based models by converting it into a probability flow ODE (Song et al., 2021b).
B.5 Vision Transformers for Diffusion Models
Diffusion transformers (Peebles & Xie, 2023) have demonstrated significant scalability and generative capabilities in image-related tasks (Esser et al., 2024; Hoogeboom et al., 2023; Hatamizadeh et al., 2024; Zhou et al., 2024). In particular, HDiT (Crowson et al., 2024) leverages the inherent hierarchical nature of visual patterns in its model design. By integrating the characteristics of Diffusion Transformer (Peebles & Xie, 2023) and Hourglass transformers (Nawrot et al., 2022), and employing local attention mechanisms (Hassani et al., 2023), HDiT offers an efficient model structure suitable for training the diffusion model in the data space.
Figure 13: Examples of problematic cases removed from the training dataset after manual curation. Low contour levels are used to highlight the heterogeneous local resolution.
Appendix C Additional Data Curation and Processing
C.1 Data Curation
Examples of the problematic cases which were excluded from the training set after manual curation are shown in Fig. 13. These includes density maps with very heterogeneous local resolutions that significant portions of the maps are poorly resolved. The EMDB IDs of the training and testing data used in this paper have been uploaded to https://figshare.com/s/9ef2614108391c04d910.
C.2 Data Standardization
We preprocessed the deposited density maps to normalize their values to a consistent range of $[0,1]$. Since the absolute values of density maps are not intrinsically meaningful and the value ranges vary significantly, we applied a uniform processing procedure as follows: Clipping: We clipped the values of each density map based on the minimum value, which is set to 00, and the $99.999$th percentile value of the density map. Scaling: Subsequently, the clipped data were scaled to the $[0,1]$ range by adjusting according to the new minimum and maximum values obtained after clipping. Normalization: we used a mean of $0.04$ and a standard deviation of $0.09$, which were calculated based on the data that have been scaled to the $[0,1]$ range.
Appendix D Additional Detailed Information of CryoFM
D.1 Implementation Details
@@
D.2 Likelihood Estimation
Given a data point ${\mathbf{x}}_{0}\in\mathbb{R}^{n}$, we estimate its likelihood by solving a probability flow ODE (Song et al., 2021b; Lipman et al., 2022; Chen et al., 2018). We start with the continuity equation of a velocity field ${\textnormal{v}}_{\Theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$:
$\frac{d}{dt}\log p_{t}({\mathbf{x}}_{t})+\nabla\cdot{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})=0.$
@@ D.2 Likelihood Estimation
$\frac{d}{dt}\left[\begin{array}[]{cc}{\mathbf{x}}_{t}\\ f_{t}\end{array}\right]=\left[\begin{array}[]{cc}{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})\\ \nabla\cdot{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})\end{array}\right],$
@@
$\frac{d}{dt}\left[\begin{array}[]{cc}{\mathbf{x}}_{t}\\ f_{t}\end{array}\right]=\left[\begin{array}[]{cc}{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})\\ {\mathbf{z}}^{\top}\nabla{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t}){\mathbf{z}}\end{array}\right],$
@@
Appendix E Additional information of the experiments
E.1 Metrics
@@
E.2 Spectral Noise Denoising
@@
Appendix F Estimation of the forward operators
F.1 Spectral noise power estimation
@@
F.2 Anisotropic noise power estimation
@@
where $h(\nu)$ is the average noise power derived from the the half-map GSFSC in Section F.1.
@@
Appendix G Additional Algorithms
@@
G.1 Missing Wedge Restoration
For certain tasks, such as missing wedge restoration, it is not necessary to compute the gradient with respect to the loss function between ${\mathbf{y}}$ and $\mathcal{A}{\mathbf{x}}$, since we can directly attain the optimum in an analytical form. The missing wedge restoration is an in-painting task, where a noiseless part of the signal is already observed. $\mathcal{A}$ is an operator extracting a part of the signal, hence the inverse of $\mathcal{A}$ just fills the observed part into the signal. In Alg. 2, we formulate it as the pseudo-inverse operator $\mathcal{A}^{\dagger}:\mathbb{R}^{m}\to\mathbb{R}^{n}$ (Kawar et al., 2021; 2022).
@@ G.1 Missing Wedge Restoration
a pretrained vector field ${\textnormal{v}}_{\Theta}:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, a forward operator $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, an obseravation ${\mathbf{y}}\in\mathbb{R}^{m}$, number of steps $N$
@@
for $t\in[1,1-\Delta t,1-2\Delta t,\cdots,\Delta t]$ do
     $\hat{{\mathbf{x}}}_{0}\leftarrow{\mathbf{x}}_{t}-t\cdot{\textnormal{v}}_{\Theta}(t,{\mathbf{x}}_{t})$
@@
     ${\mathbf{x}}_{t-\Delta t}\leftarrow\frac{t-\Delta t}{t}\cdot{\mathbf{x}}_{t}+\frac{\Delta t}{t}\cdot\hat{{\mathbf{x}}}^{\prime}_{0}$
@@
G.2 Ab initio Modeling
For ab initio modeling, the $k$-th observed volume (we will name it the projection later) is $\mathcal{A}^{(k)}({\bm{V}})=\mathcal{P}(\phi^{(k)},{\bm{V}})$, where $\phi^{(k)}\in\mathbb{SO}(3)\times\mathbb{R}^{2}$ is the pose. The posterior sampling of ab initio modeling is more complex than the previous tasks in that the pose $\phi^{(k)}$ can not be easily pre-determined. The setting is closely related to blind diffusion posterior sampling (Chung et al., 2023a; Kapon et al., 2024; Laroche et al., 2024; Gan et al., 2024), where the forward operator is unknown. We find $\phi^{(k)}$ in the sampling process, by minimizing the discrepancy between the projection and the observation ${\bm{y}}^{(k)}$ using a coarse-to-fine strategy (Zhong et al., 2021b). The discrepancy is defined by the negative correlation between a projection ${\mathbf{p}}$ and an observation ${\mathbf{y}}$: